# HIF: The hypergraph interchange format for higher-order networks

### v1

@@ Abstract
Many empirical systems contain complex interactions of arbitrary size, representing, for example, chemical reactions, social groups, co-authorship relationships, and ecological dependencies. These interactions are known as higher-order interactions and the collection of these interactions comprise a higher-order network, or hypergraph. Hypergraphs have established themselves as a popular and versatile mathematical representation of such systems and a number of software packages written in various programming languages have been designed to analyze these networks. However, the ecosystem of higher-order network analysis software is fragmented due to specialization of each software’s programming interface and compatible data representations. To enable seamless data exchange between higher-order network analysis software packages, we introduce the Hypergraph Interchange Format (HIF), a standardized format for storing higher-order network data. HIF supports multiple types of higher-order networks, including undirected hypergraphs, directed hypergraphs, and simplicial complexes, while actively exploring extensions to represent multiplex hypergraphs, temporal hypergraphs, and ordered hypergraphs. To accommodate the wide variety of metadata used in different contexts, HIF also includes support for attributes associated with nodes, edges, and incidences. This initiative is a collaborative effort involving authors, maintainers, and contributors from prominent hypergraph software packages. This project introduces a JSON schema with corresponding documentation and unit tests, example HIF-compliant datasets, and tutorials demonstrating the use of HIF with several popular higher-order network analysis software packages.
@@
complex network data; higher-order network science; hypergraph; directed hypergraph; simplicial complex; dataset standards
I Background
@@
The growth of network science as a field has been fueled by the availability of large-scale datasets from a variety of sources including the internet [1], co-authorship networks [2, 3], global-scale transportation [4], social interactions [5, 6, 7], biological processes [8, 9], social media [10, 11, 12], and brain connectivity [13], among others. As the plethora of data has grown, standardized formats for pairwise network datasets have been developed such as GraphML [14] and Pajek [15] as well as network dataset repositories such as SNAP [16], Netzschleuder [17], ICON [18], and Konect [19]. These advances, however, have been developed with pairwise networks primarily in mind, which solely model dyadic interactions.
@@
Higher-order networks generalize the notion of pairwise networks, representing interactions of arbitrary size. These networks can more naturally model certain empirical systems such as co-authorship networks, chemical reactions, human social networks, and biological interactions, which abound in multi-way interactions. This can be useful for studying nonlinear dynamical processes [20], the evolution of groups [21], and protein-protein interactions [22], for example. There are many robust software packages for storing, analyzing, and visualizing higher-order networks [23, 24, 25, 26, 27, 28, 29, 30, 31], but no corresponding data standards for facilitating cross-package compatibility. In this paper, co-authored by leaders and representatives of a number of these software packages, we explain the need for such a standard and the context in which it fits; introduce the Hypergraph Interchange Format (HIF) as a data standard for higher-order networks; describe integration with existing higher-order network software libraries; and lastly, demonstrate how it can be used to unlock the strengths of different higher-order network software packages.
I.1 Existing higher-order network data
@@
Higher-order network data remains relatively scattered, with multiple versions of the same datasets often distributed across several repositories and lacking a consistent format. For example, the webpage curated by Austin Benson [32] features many datasets, primarily from the computer science domain, but is no longer actively maintained. There exist several actively maintained data repositories: XGI-DATA [31], which hosts many datasets on Zenodo  [^1] See https://zenodo.org/communities/xgi. and is integrated with the XGI software API; HypergraphRepository [34], which offers a collection of empirical hypergraph datasets along with a web application for filtering and downloading them; and the Hypergraphx [26] software library, which is accompanied by a diverse collection of datasets spanning multiple domains and hypergraph types [35]. Additionally, higher-order network datasets can sometimes be found in traditional network repositories such as SNAP [16], ICON [18], and Netzschleuder [17], typically represented as bipartite graphs. Notably, there may be repetitions and overlapping datasets across these repositories, resulting in a redundant and fragile ecosystem. Furthermore, each repository currently uses its own proprietary format, which is not directly interoperable with other software libraries.
@@
HIF addresses this challenge by offering a consistent, expressive, and extensible JSON-based schema for encoding both the topology and metadata of hypergraphs. Its language-agnostic design enables seamless exchange between libraries implemented in Python, C++, Julia, JavaScript, and other environments, promoting modular, interoperable workflows while reducing redundant engineering effort. HIF serves as a unifying layer for the higher-order network science ecosystem, allowing software tools and pipelines to communicate through a shared representation. This facilitates reproducible research and fosters a more cohesive, collaborative development environment, similar to the role of the Open Neural Network Exchange (ONNX) format in machine learning [36] and GraphML in traditional graph processing [14].
II Higher-Order Networks
@@
The term ”higher-order network” has become associated with a variety of data objects related to hypergraphs. The literature on hypergraphs is extensive, and there are multiple ways to define them; hypergraph representations must specify how to handle issues like the role of duplicate hyperedges, self-loops, empty hyperedges, isolated vertices, and metadata. We refer the reader to two classic publications about hypergraphs [37, 38] as well as more recent treatments of the larger space of higher-order networks [39, 40, 41]. And while the foundational mathematics of hypergraphs continues to evolve, including in a category-theoretical framework [42], our approach is to seek an appropriate general grounding sufficient to accommodate the widest range of both current hypergraph modeling approaches and current data science needs, while avoiding some of the more subtle complexities which can arise. To that end, we now introduce just a few concepts of the most direct relevance to HIF formally, while noting both similarities and differences with some of the most widely circulated concepts as used in the literature.
@@
For our purposes, a hypergraph is a system $H=(V,E,I)$ where $V=\{v\}$ is a finite, non-empty set of vertices or nodes, $E=\{e\}$ is a finite, non-empty set of edges or hyperedges, and $I\subseteq E\times V$ is a set of incidences, that is, pairs $(v,e)$ of nodes and edges.
@@
It may be valuable to understand how this definition may differ from some readers’ expectations. First, note that an edge $e\in E$ can be mapped to the collection of vertices with which it has an incidence: $e\rightarrow\{v\in V:(v,e)\in E\}$. With this in mind, we will also feel free to simply think of an edge as a subset of vertices, and notate $e\subseteq V$, thereby recovering the traditional sense of a hyperedge as a collection of vertices of arbitrary size. And indeed, we could also notate $H=(V,E)$ and simply understand that each edge $e\in E$ is a subset $e\subseteq V$. But what’s critical to note is that then $E$ is not actually a set, but rather, using our definition $E$ is an ”indexed family” of hyperedges possibly containing duplicates: there can be two distinct edges $e,f\in E$ which map to the same set of vertices, so that $e=f$. So while mathematically, our structure may be properly called a multi-hypergraph, for this paper, we will simply call these all hypergraphs, while still keeping this issue firmly in mind.
@@
Rather more well understood is how any particular edge is a subset $e\subseteq V$ of arbitrary size. If $e$ maps to only two vertices, then that $e$ is a traditional (undirected) graph edge; and indeed, if this is true for all edges $e\in E$, then $H$ is called a 2-uniform hypergraph, which is just a graph: all graphs are hypergraphs. But $e$ may also have more than two nodes (the traditional sense of ”higher-order”), one node (singleton hyperedges), or even none (empty hyperedges are permitted).
@@
Another prominent structure in higher-order network theory is that of an abstract simplicial complex (ASC). An ASC is a hypergraph, but one which is ”downward closed” in that for every hyperedge $e\in E$, then all nonempty collections of vertices $f\subseteq e$ contained in $e$ are also in the hypergraph: $f\in E$.
@@
Essential to HIF is that it supports attributed hypergraphs: nodes $v\in V$ and edges $e\in E$ have arbitrary properties. But of great significance is that in addition, HIF also supports properties on incidences $(v,e)\in I$. In other words, a node $v\in V$ can have one property when associated with an incidence $(v,e)$ involving one edge $e\in E$ in which it is contained ($v\in e$), but could have quite a different property when associated with a different incidence $(v,f)$ for a different edge $f\in E$, for which also $v\in f$.
Incidence properties are essential for supporting critical features like ”edge-dependent weights” [43] used in, e.g., random walk models on hypergraphs [44]. But another critical use of incidence properties in HIF is to represent directed hypergraphs [45, 46]. Directed hypergraphs comprise directed hyperedges each of which is an ordered pair of sets: the ”tail”, specifying the inputs in the multi-way interaction, and the ”head”, specifying the outputs. In contrast to prior literature requiring that the head and tail be disjoint [47], here we generalize this framework and allow nodes to belong to both the tail and head. An incidence property can be used to indicate whether, for any particular hyperedge $e\in E$, a vertex $v\in e$ is in the tail, the head, or both.
III The Hypergraph Interchange Format
@@
The Hypergraph Interchange Format is a standard for storing higher-order networks in JavaScript Object Notation (JSON) format. JSON is chosen for its interpretability; while there are more efficient formats such as Parquet, CSV, GraphQL, etc., these data structures either lack human-readability or support for the rich attributes often accompanying higher-order data. JSON is supported by all major programming languages and JSON validation libraries written in Python, R, and Julia provide a user-friendly way to verify datasets, as seen in Fig. 2, making this a suitably popular file format. HIF offers support not only for network-level metadata, but also for attributes corresponding to nodes, hyperedges, and incidences and it does so in an interpretable way. This specification also handles several different types of higher-order networks and is flexible enough to accommodate additional higher-order representations in the future.
@@
III.1 The Hypergraph Interchange Format specification
Figure 1: The hierarchy of properties in the HIF JSON Schema.
@@ III.1 The Hypergraph Interchange Format specification
network-type (optional) enum ”directed”, ”undirected”, or ”asc”
@@
direction (optional) enum ”head” or ”tail”
@@
The HIF specification organizes data about a higher-order network in five different categories via five top-level fields in the JSON file: network-type, metadata, nodes, edges, and incidences. The text corresponding to the network-type field indicates the type of the higher-order network and offers additional guidance for processing the dataset which we will discuss in detail below. The data corresponding to the metadata field is a dictionary-like object representing network-level attributes such as the name of the dataset, when it was collected, the author of the dataset, corresponding references, and other high-level information enabling greater dataset interpretability. The nodes and edges fields store data records for the nodes and edges, respectively. These records can be used to specify metadata corresponding to each node and hyperedge, or as we will discuss in more detail below, to specify empty edges or isolated nodes. Lastly, the incidences field stores data records detailing node-hyperedge relationships (known as an incidence). Below we detail more of the structure as well as some of the implications of this specification.
@@
We start with the incidences data, because this is the only top-level attribute which explicitly encodes the complex relationships between nodes via hyperedges. The incidences field is the only required field in HIF-compliant files, and a file only including this field corresponds to a higher-order network without any network, node, or hyperedge attributes. The incidence data are composed of a list of entries known as records, where each record indicates a relationship between a node and a hyperedge, known as an incidence. For example, {”edge”: 1, ”node”: 3} indicates that node 3 belongs to hyperedge 1. The collection of all records where ”edge” is 1 forms a hyperedge, e.g., the collection of records {”edge”: 1, ”node”: 3}, {”edge”: 1, ”node”: 4},{”edge”: 1, ”node”: 7} corresponds to a hyperedge comprising the interaction between nodes 3, 4, and 7. Additional structural features which can be specified are weight and, in the case of directed hypergraphs, direction. These properties are added as top-level attributes in each record; e.g., {”edge”: 1, ”node”: 3, ”weight”: 2} specifies an incidence with an edge-dependent node weight of 2 and {”edge”: 1, ”node”: 3, ”direction”: ”tail”} indicates that node 3 is in the tail of directed hyperedge 1. Incidences can also have properties, which indicate hyperedge-dependent nodal attributes. All of these attributes are bundled as dictionary-like objects in each incidence record under the attrs field, e.g., {”edge”: 1, ”node”: 3, ”attrs”: {”role”: ”PI”}} might indicate node 3’s role as a PI on project 1.
@@
Similarly, we can specify attributes for nodes and hyperedges by creating records corresponding to the nodes and edges fields, respectively. For example, {”node”: 0, ”attrs”: {”height”: 176, ”weight”: 143}} corresponds to node 0 with a height of 176 and weight of 143, and {”edge”: 1, ”attrs”: {”duration”: 93, ”setting”: ”coffee shop”}} corresponds to an 93 minute interaction at a coffee shop. Another use of these fields is to specify empty hyperedges and isolated nodes which don’t participate in these incidence relationships and thus are not listed in the incidence records. We can add these nodes and edges as node and edge records with or without attributes to ensure that they are present in the higher-order network.
@@
The network-type field can take one of three values: ”asc”, ”directed”, and ”undirected”. This field offers additional guidance for processing the higher-order network. For example, using the ”asc” keyword indicates that, even if all subfaces are not specified, that the network should be downward closed. The best practice for storing a simplicial complex is to only store the maximal faces and subfaces with attributes to minimize the storage required, but this is a guideline, not a hard requirement. When using the ”directed” keyword, this indicates that the library should expect the ”direction” field indicating whether the incidence forms part of the head or tail of the directed hyperedge. When the network-type is not specified, it is assumed that the type is an undirected hypergraph.
III.2 Checking HIF compliance

### v2

@@ Abstract
Many empirical systems contain complex interactions of arbitrary size, representing, for example, chemical reactions, social groups, co-authorship relationships, and ecological dependencies. These interactions are known as higher-order interactions and the collection of these interactions comprise a higher-order network, or hypergraph. Hypergraphs have established themselves as a popular and versatile mathematical representation of such systems and a number of software packages written in various programming languages have been designed to analyze these networks. However, the ecosystem of higher-order network analysis software is fragmented due to specialization of each software’s programming interface and compatible data representations. To enable seamless data exchange between higher-order network analysis software packages, we introduce the Hypergraph Interchange Format (HIF), a standardized format for storing higher-order network data. HIF supports multiple types of higher-order networks, including undirected hypergraphs, directed hypergraphs, and abstract simplicial complexes, while actively exploring extensions to represent multiplex hypergraphs, temporal hypergraphs, and ordered hypergraphs. To accommodate the wide variety of metadata used in different contexts, HIF also includes support for attributes associated with nodes, edges, and incidences. This initiative is a collaborative effort involving authors, maintainers, and contributors from prominent hypergraph software packages. This project introduces a JSON schema with corresponding documentation and unit tests, example HIF-compliant datasets, and tutorials demonstrating the use of HIF with several popular higher-order network analysis software packages.
keywords:
@@
network, higher-order, hypergraph, directed hypergraph, simplicial complex, data standard
[…]
1 Introduction
@@
The growth of network science as a field has been fueled by the availability of large-scale datasets from a variety of sources including the internet albert_diameter_1999, co-authorship networks barabasi_evolution_2002; newman_structure_2001, global-scale transportation brockmann_hidden_2013, social interactions cattuto_dynamics_2010; stehle_highresolution_2011; adams_gathering_2019, biological processes barabasi_network_2004; jeong_largescale_2000, social media jackson_hashtagactivism_2020; morstatter_sample_2013; ugander_structural_2012, and brain connectivity sporns_organization_2004, among others. As the plethora of data has grown, standardized formats for pairwise network datasets have been developed such as GraphML brandes_graph_2010 and Pajek nooy_exploratory_2018 as well as network dataset repositories such as SNAP leskovec_snap_2014, Netzschleuder peixoto_netzschleuder_2021, ICON clauset_colorado_2016, and Konect kunegis_konect_2013. These advances, however, have been developed with networks primarily in mind, which solely model dyadic interactions.
@@
Higher-order networks generalize the notion of networks, representing interactions of arbitrary size. These networks can more naturally model certain empirical systems such as co-authorship networks, chemical reactions, human social networks, and biological interactions, which abound in multi-way interactions. This can be useful for studying nonlinear dynamical processes neuhauser_multibody_2020, the evolution of groups iacopini_temporal_2024, and protein-protein interactions murgas_hypergraph_2022, for example. There are many robust software packages for storing, analyzing, and visualizing higher-order networks failla_attributed_2023; pickard_hat_2023; diaz_hypergraphsjl_2022; lotito_hypergraphx_2023; praggastis_hypernetx_2024; badie-modiri_reticula_2023; spagnuolo_analyzing_2020; hajij_topox_2024; landry_xgi_2023, but no corresponding data standards for facilitating cross-package compatibility. In this paper, co-authored by leaders and representatives of a number of these software packages, we explain the need for such a standard and the context in which it fits; introduce the Hypergraph Interchange Format (HIF) as a data standard for higher-order networks; describe integration with existing higher-order network software libraries; and lastly, demonstrate how it can be used to unlock the strengths of different higher-order network software packages.
1.1 Existing higher-order network data
@@
Higher-order network data remains relatively scattered, with multiple versions of the same datasets often distributed across several repositories and lacking a consistent format. For example, the webpage curated by Austin Benson benson_data_2021 features many datasets, primarily from the computer science domain, but is no longer actively maintained. There exist several actively maintained data repositories: XGI-DATA landry_xgi_2023, which hosts many datasets on Zenodo  [^1] See https://zenodo.org/communities/xgi. and is integrated with the XGI software API; HypergraphRepository antelmi_hypergraphrepository_2024, which offers a collection of empirical hypergraph datasets along with a web application for filtering and downloading them; and the Hypergraphx lotito_hypergraphx_2023 software library, which is accompanied by a diverse collection of datasets spanning multiple domains and hypergraph types HGXTeamData. Additionally, higher-order network datasets can sometimes be found in traditional network repositories such as SNAP leskovec_snap_2014, ICON clauset_colorado_2016, and Netzschleuder peixoto_netzschleuder_2021, typically represented as bipartite graphs (see discussion in Section 2 below). Notably, there may be repetitions and overlapping datasets across these repositories, resulting in a redundant and fragile ecosystem. Furthermore, each repository currently uses its own proprietary format, which is not directly interoperable with other software libraries.
@@ 1.2 Contributions
HIF addresses this challenge by offering a consistent, expressive, and extensible JSON-based schema for encoding both the topology and metadata of hypergraphs. Its language-agnostic design enables seamless exchange between libraries implemented in Python, C++, Julia, JavaScript, and other environments, promoting modular, interoperable workflows while reducing redundant engineering effort. HIF serves as a unifying layer for the higher-order network science ecosystem, allowing software tools and pipelines to communicate through a shared representation. This facilitates reproducible research and fosters a more cohesive, collaborative development environment, similar to the role of the Open Neural Network Exchange (ONNX) format in machine learning onnx2024 and GraphML in traditional graph processing brandes_graph_2010.
2 Higher-Order Networks as Hypergraphs
Network science is grounded on the mathematical representation of networks as undirected or directed graphs. Similarly the term “higher-order network” has become associated with a variety of data objects related to undirected and directed hypergraphs. The literature on hypergraphs is extensive, and we refer the reader to two classic publications about hypergraphs berge_hypergraphs_1984; bretto_hypergraph_2013 as well as more recent treatments of the larger space of higher-order networks battiston_networks_2020; torres_why_2021a; joslyn_hypernetwork_2021; bick_what_2023a.
@@ 2 Higher-Order Networks as Hypergraphs
The foundational mathematics of hypergraphs continues to evolve, including in a category-theoretical framework grilliette_incidence_2023, producing multiple ways to axiomatize hypergraph formalisms. Our approach for this paper is to seek an appropriate general grounding sufficient to accommodate the widest range of both current hypergraph modeling approaches and current data science needs, while avoiding some of the more subtle complexities which can arise especially around edge cases like empty and duplicate hyperedges, isolated and redundant vertices, empty hypergraphs, self-loops, and related things. To that end, we now introduce just a few concepts of the most direct relevance to HIF formally, while noting both similarities and differences with some of the most widely circulated concepts as used in the literature.
@@
Thus for our purposes, a hypergraph is a system $H=(V,E,I)$ where $V=\{v\}$ is a finite, non-empty set of vertices or nodes, $E=\{e\}$ is a finite, non-empty set of edges or hyperedges, [^2] We will commonly refer to “edges” to mean hyperedges for convenience and when clear from context. and $I\subseteq V\times E$ is a set of incidences, that is, pairs $(v,e)$ of nodes and edges.
@@
It may be valuable to understand how this definition may differ from some readers’ expectations. First, note that an edge $e\in E$ can be mapped to the collection of vertices with which it has an incidence: $e\mapsto\{v\in V:(v,e)\in I\}$. With this in mind, we will also feel free to simply think of an edge as a subset of vertices, and notate $e\subseteq V$, thereby recovering the traditional sense of a hyperedge as a collection of vertices of arbitrary size. And indeed, we could also notate $H=(V,E)$ and simply understand that each edge $e\in E$ is a subset $e\subseteq V$. But what’s critical to note is that then this $E$ is not actually a set, but rather, using our definition $E$ would be a multiset or bag of hyperedges, and so possibly containing duplicates: there can be two distinct edges $e,f\in E$ which map to the same set of vertices, so that $e=f$. So while mathematically, our structure may be properly called a multi-hypergraph, for this paper, we will simply call these all hypergraphs, while still keeping this issue firmly in mind.
@@
Rather more well understood is how any particular edge is a subset $e\subseteq V$ of arbitrary size. If $e$ maps to only two vertices, then that $e$ is a traditional (undirected) graph edge; and indeed, if this is true for all edges $e\in E$, then $H$ is called a 2-uniform hypergraph, which is just a graph: all graphs are hypergraphs. But $e$ may also have more than two nodes (the traditional sense of “higher-order”), one node (singleton hyperedges), or even none (empty hyperedges are permitted).
HIF supports directed hypergraphs ausiello_directed_2001; gallo_directed_1993a, which comprise directed hyperedges $e\in E$. Each directed hyperedge is an ordered pair $e=(t,h)$ of subsets of vertices: the “tail” $t\subseteq e$ specifies the inputs in the multi-way interaction, and the “head” $h\subseteq e$, specifies the outputs. While the tail and head cover the hyperedge in that $t\cup h=e$, in contrast to some researchers  jost_hypergraph_2019 who require that the head and tail be disjoint, here we generalize this framework and allow nodes to belong to both the tail and head.
@@
Another prominent structure in higher-order network theory is that of an abstract simplicial complex (ASC). An ASC is a hypergraph, but one which is “downward closed” in that for every hyperedge $e\in E$, then all nonempty collections of vertices $f\subseteq e$ contained in $e$ are also in the hypergraph: $f\in E$. ASCs are used extensively in higher-order network science (joslyn_hypernetwork_2021), for example when we seek to analyze the topological properties of hypergraphs mathematically as structures with multiple multi-dimensional components glued together, or when semantically it is understood in some application that when a relationship holds for a set of entities, it also holds for all its subsets.
@@
Essential to HIF is that it supports attributed hypergraphs: nodes $v\in V$ and edges $e\in E$ can have arbitrary properties. But of great significance is that in addition, HIF also supports properties on the incidences $(v,e)\in I$. In other words, a node $v\in V$ can have one property when associated with an incidence $(v,e)$ involving one edge $e\in E$ in which it is contained ($v\in e$), but could have quite a different property when associated with a different incidence $(v,f)$ for a different edge $f\in E$, for which also $v\in f$. An incidence property can be used to represent directed hyperedges, in that for any particular hyperedge $e\in E$ and vertex $v\in e$, an incidence property can be used to encode whether $v$ is in the tail $v\in t$, the head $v\in h$, or in both. Incidence properties are also essential for supporting critical features like “edge-dependent weights” chitra_random_2019 used in, e.g., random walk models on hypergraphs hayashi_hypergraph_2020.
It is also worth noting a few additional properties of hypergraphs and mathematical structures related to hypergraphs. While these aren’t represented in HIF proper, they are important for applications and understanding, and are mentioned elsewhere in this paper in the context of their implementation in some of the hypergraph software systems supporting HIF.
First, the incidence relation $I\subseteq E\times V$ is commonly represented as a Boolean incidence matrix, sometimes also notated (perhaps abusively) as just $I$, with $|V|$ rows and $|E|$ columns, where $I[i,j]=1$ when $(v,e)\in I$ and 0 when $(v,e)\not\in I$. Matrix operations on $I$ can be very useful in applications, as we will see below. Additionally, every hypergraph $H=(V,E,I)$ has an equivalent dual form $H^{*}=(E,V,I^{-1})$ where nodes and edges are swapped, and $I^{-1}$ is the inverse of the incidence relation, yielding the dual incidence matrix $I^{T}$ as the transpose.
Every hypergraph $H$ is also equivalently represented as a bipartite graph $B=(V\sqcup E,A)$ on a new set of vertices as the disjoint union of the nodes $V$ and hyperedges $E$ of the original hypergraph, and with a new set of graph edges $A\subseteq\binom{V\sqcup E}{2}$, but with the bipartite condition that for all edges $\{x,y\}\in A$, exactly one of $x$ or $y$ are in $V$, with the other in $E$. Undirected hypergraphs yield undirected bipartite graphs; directed hypergraphs yield directed bipartite graphs; and vice versa in both cases. While hypergraphs are bijective with bipartite graphs, they can represent different concepts and functions, and can be amenable to different algorithms: hypergraphs are about connectivity and path following, where bipartite graphs are about combinatoric and matching questions. And from a mathematical perspective, they may also have different transformations which yield them in different categories (grilliette_incidence_2023).
Hypergraphs operate in a broader mathematical ecosystem which also includes graphs. We have already seen how every graph is a 2-uniform hypergraph. Additionally, given a hypergraph $H=(V,E,I)$, it is common to construct a graph $(V,C)$ on its vertex set, called a clique expansion, 2-section, or underlying graph. Here $C$ is the union of the complete graphs (cliques) formed on each hyperedge $e\in E$. Thereby $C$ consists of all the pairs of vertices (graph edges on $V$) included in some hyperedge, such that for all $\{u,v\}\in C,\exists e\in E,\{u,v\}\subseteq e$. Note that the adjacency matrix of the 2-section is given by $I\times I^{T}$ as a matrix operation. Finally, it is also common to build the line graph of a hypergraph $H=(V,E,I)$ as a graph $(E,L)$ on the edges taken as vertices, where a pair of edges $\{e,f\}$ is included in $L$ if they intersect: $L=\{e,f\in E:e\cap f\neq\emptyset\}$. The line graph structure allows reasoning about the relationships between the hyperedges, as opposed to the vertices, and is thus closely related to the dual hypergraph $H^{*}$: indeed, the line graph of $H$ is just the 2-section of the dual $H^{*}$, and vice versa, where now the adjacency matrix of the line graph is $I^{T}\times I$.
3 The Hypergraph Interchange Format
@@
The Hypergraph Interchange Format is a standard for storing higher-order networks in JavaScript Object Notation (JSON) format. JSON is chosen for its interpretability; while there are more efficient formats such as Parquet, CSV, GraphQL, etc., these data structures either lack human-readability or support for the rich attributes often accompanying higher-order data. JSON is supported by all major programming languages, and JSON validation libraries written in Python, R, and Julia provide a user-friendly way to verify datasets, as seen in Figure 2, making this a suitably popular file format. HIF offers support not only for network-level metadata, but also for attributes corresponding to nodes, hyperedges, and incidences, and it does so in an interpretable way. This specification also handles several different types of higher-order networks and is flexible enough to accommodate additional higher-order representations in the future.
@@
3.1 The Hypergraph Interchange Format specification
(a) The hierarchy of properties in the HIF JSON Schema.
"network-type": "undirected",
"metadata": {
"title": "HIF publications dataset",
"collection-date": "2025-07-04",
"doi": "10.5281/zenodo.15802760"
"incidences": [
{"edge": 10, "node": 20, "weight": 2},
{"edge": 30, "node": 20, "weight": 3}
"nodes": [
{"node": 20, "attrs":
{"institutions": ["PNNL", "UVA"]}}
"edges": [
{"edge": 10, "attrs":
{"source": "Arxiv",
"tags": ["Machine Learning", "Artificial Intelligence"]}},
{"edge": 30, "attrs":
{"source": "DBLP", "tags": ["Hypergraph"]}}
(b) An example HIF document in JSON format.
Figure 1: The JSON Schema is depicted as a graph, showing the hierarchy of properties for each data record. The example demonstrates the use of most properties including metadata for interpretability.
@@ 3.1 The Hypergraph Interchange Format specification
network-type (optional) enum “directed”, “undirected”, or “asc”
@@
direction (optional) enum “head” or “tail”
@@
The HIF specification organizes data about a higher-order network in five different categories via five top-level fields in the JSON file: network-type, metadata, nodes, edges, and incidences as can be seen in Table 1. The text corresponding to the network-type field indicates the type of the higher-order network and offers additional guidance for processing the dataset which we will discuss in detail below. The data corresponding to the metadata field is a dictionary-like object representing network-level attributes such as the name of the dataset, when it was collected, the author of the dataset, corresponding references, and other high-level information enabling greater dataset interpretability. The nodes and edges fields store data records for the nodes and edges, respectively. These records can be used to specify metadata corresponding to each node and hyperedge, or as we will discuss in more detail below, to specify empty edges or isolated nodes. Lastly, the incidences field stores data records detailing node-hyperedge relationships (known as an incidence). Below we detail more of the structure as well as some of the implications of this specification.
@@
We start with the incidences data, because this is the only top-level attribute which explicitly encodes the complex relationships between nodes via hyperedges. The incidences field is the only required field in HIF-compliant files, and a file only including this field corresponds to a higher-order network without any network, node, or hyperedge attributes. The incidence data are composed of a list of entries known as records, where each record indicates a relationship between a node and a hyperedge, known as an incidence. For example, {“edge”: 1, “node”: 3} indicates that node 3 belongs to hyperedge 1. The collection of all records where “edge” is 1 forms a hyperedge, e.g., the collection of records {“edge”: 1, “node”: 3}, {“edge”: 1, “node”: 4}, {“edge”: 1, “node”: 7} corresponds to a hyperedge comprising the interaction between nodes 3, 4, and 7. Additional structural features which can be specified are weight and, in the case of directed hypergraphs, direction. These properties are added as top-level attributes in each record; e.g., {“edge”: 1, “node”: 3, “weight”: 2} specifies an incidence with an edge-dependent node weight of 2 and {“edge”: 1, “node”: 3, “direction”: “tail”} indicates that node 3 is in the tail of directed hyperedge 1. Incidences can also have properties, which indicate hyperedge-dependent nodal attributes. All of these attributes are bundled as dictionary-like objects in each incidence record under the attrs field, e.g., {“edge”: 1, “node”: 3, “attrs”: {“role”: “PI”}} might indicate node 3’s role as a PI on project 1.
@@
Similarly, we can specify attributes for nodes and hyperedges by creating records corresponding to the nodes and edges fields, respectively. For example, {“node”: 0, “attrs”: {“height”: 176, “weight”: 143}} corresponds to node 0 with a height of 176 and weight of 143, and {“edge”: 1, “attrs”: {“duration”: 93, “setting”: “coffee shop”}} corresponds to an 93 minute interaction at a coffee shop. Another use of these fields is to specify empty hyperedges and isolated nodes which don’t participate in these incidence relationships and thus are not listed in the incidence records. We can add these nodes and edges as node and edge records with or without attributes to ensure that they are present in the higher-order network.
@@
The network-type field can take one of three values: “asc”, “directed”, and “undirected”. This field offers additional guidance for processing the higher-order network. For example, using the “asc” keyword indicates that, even if all subfaces are not specified, that the network should be downward closed. The best practice for storing an abstract simplicial complex is to only store the maximal faces and subfaces with attributes to minimize the storage required, but this is a guideline, not a hard requirement. When using the “directed” keyword, this indicates that the library should expect the “direction” field indicating whether the incidence forms part of the head or tail of the directed hyperedge. When the network-type is not specified, it is assumed that the type is an undirected hypergraph.
The metadata field can take arbitrary information to support the dataset under study. The schema allows any object in the metadata property, including deeply nested structures. Dataset authors are encouraged to use a flat object and to use only lowercase characters and hyphens, a convention commonly known as “dash case”. This can be seen in Figure 1 where the “Creation Date” property name is encoded as “creation-date”, and “DOI” is encoded as “doi”.
3.2 Checking HIF compliance